Chester Corp. is downsizing the size of their workforce by 10% (to the nearest person) next year from various strategic initiatives.   How much will the company pay in separation costs if each worker receives $5,000 when separated?

complement needed 587

a.$295,000
b.$2,640,00
c.$1,056,00
d.$118,000
Answer #1

1. Initial workforce = 587

New workforce = 587*0.9 = 528.3 or 528

Difference = 587 - 528 = 59

Money received by each worker = $5000

Total money paid by company = 5000*59 = $295,000

Hence total amount paid by company in separation cost = $295,000